OpenStudy (anonymous):

if u turn 1.275 into a percent. would it be .44

OpenStudy (saifoo.khan):

1.275 means when u are going to solve the decimal it will be like, \[\frac{1275}{1000}\] And now u already know how to convert a number to a percentage, \[\frac{1275}{1000} \times 100\]

OpenStudy (saifoo.khan):

So now, \[\frac{1275}{10\cancel 0\cancel 0} \times 1\cancel 0 \cancel 0\]

OpenStudy (saifoo.khan):

Finally, \[\frac{1275}{10}\]Move the deciaml place to the left and u are done! :D 127.5 % Answer

OpenStudy (anonymous):

Thanks

OpenStudy (saifoo.khan):

Please hit good answer.

OpenStudy (anonymous):

ok